An oxide of nitrogen contains 63.1% oxygen and has a molar mass of 76.0 g/mol.







a. What is the empirical formula of the compound?







b. What is the molecular formula of the compound?

1
Expert's answer
2022-02-24T00:57:02-0500

% of nitrogen=100-63.1=36.9

Moles of nitrogen= 36.9÷14 =2.63571

Moles of oxygen = 63.1÷ 16= 3.94375

Mole ratio, N:O = (2.63571÷2.63572):(3.94375÷2.63572)

= 1:1.5

=2:3

The empirical formula is N2O3


Empirical formula mass= 14×2+16×3=76

Empirical formula mass=molecular formula mass, hence the two formulas are the same.

Molecular formula is N2O3O